A point mass is oscillating due to a restoring force. The amplitude of its oscillation is 0.2 m and the force constant is 100 Nm−1 . What is the total energy of the system?
1
Expert's answer
2019-08-14T09:51:01-0400

The total energy of the spring-mass system

"E=\frac{mv^2}{2}+\frac{kx^2}{2}=\frac{kx_{max}^2}{2}"

"=\frac{100\:\rm{N/m}\times (0.2\:\rm{m})^2}{2}=2\:\rm{J}"